# From time crystals to time quasicrystals: Exploring novel phases in transverse field Ising chains

### v1

@@ III.1 Stability of the TC phase under interactions
[…]
@@
In Fig. 2, we present the time evolution of the magnetization $m^{z}$ in the ferromagnetic many-body localized (FM-MBL) phase for varying interaction strengths. The observed time dependence of $m^{z}$ reveals significant oscillatory behavior, characterized by a period of $2T$, superimposed on an oscillatory-decaying envelope that is dependent on the interaction strength $\lambda$. Specifically, for a finite-size system, the envelope exhibits a prolonged time period at low interaction strengths, and diminishes as $\lambda$ increases. Furthermore, we have computed the Fourier transform of $m^{z}$ as a function of frequency for various $\lambda$ values. For small values of $\lambda$, during the observed time window ($t/T\in[0,160]$), the magnetization envelope curve remains smooth without any knots, and the FFT reveals a distinct subharmonic peak at $\omega=\pi$. As $\lambda$ increases, this peak begins to split due to the interference of magnetization oscillations and eventually disappears, indicating the complete destruction of the TC phase. Extending the time window for analyzing the system’s magnetization increases the probability of knot formation, leading to a double peak in the FFT even for small interaction values.
@@
The observed oscillatory-decaying envelope in the magnetization can be attributed to finite-size effects. To substantiate this assertion, we present in Fig. 3 the time evolution of the magnetization $m^{z}$ at a coupling strength $\lambda<J$ for chains of varying lengths. Notably, as the chain length increases, the period of the envelope oscillations also increases, suggesting that in the limit of infinite chain length ($L\rightarrow\infty$), the envelope disappears, leaving only stable oscillations with a period of $2T$. This behavior indicates the existence of a robust TC for interaction strengths below a critical value $\lambda_{c}$, with $\lambda_{c}$ being dependent on the system size. Furthermore, in the inset of Fig. 3, we present the FFT of $m^{z}$ for different system lengths. For smaller systems, the FFT reveals a double peak structure, reflecting the beating of magnetization oscillations. As the system size increases, these double peaks converge into a single peak at frequency $\omega=\pi$. Continued increases in system size result in a sharper peak that becomes increasingly concentrated around $\omega=\pi$.
[…]
@@
We can also evaluate the stability of the TC phase against interactions by examining how the absolute value of the magnetization, $|m^{z}|$, varies over time. As shown in Fig. 4, $|m^{z}|$ remains constant over time for interaction strengths below a critical value $\lambda_{c}$, indicating that the TC phase is stable in this regime. The critical value $\lambda_{c}$ is dependent on the system size; for a chain of length $L=12$, it is approximately $\lambda_{c}\simeq 0.4$.
Figure 4: The absolute value of the magnetization $m^{z}$ for a chain of length $L=12$ for different values of $\lambda$, with a coupling constant $J=5.5$ and an external field $h=0.3$. The TC remains stable for interaction strengths that are weaker than $\lambda_{c}\simeq 0.4$.
III.2 Stability of the TC phase under imperfect rotations
@@
We now examine the stability of the TC phase observed in the ITF model under conditions of imperfect rotations. Instead of executing a perfect spin-flip rotation (i.e., a rotation around the $x$-axis by an angle of $\frac{\pi}{2}$), we implement an imperfect rotation characterized by an angle $\theta=\frac{\pi}{2}-\epsilon$, where $\epsilon$ is a small positive value ($\epsilon<1$). Consequently, the time evolution operator is expressed as in Eq. (2) with the time parameters set to $t_{2}=1$ and $t_{1}=\frac{\pi}{2}-\epsilon$.
@@
The time-dependent behavior of the magnetization along the $z$-direction is depicted in Fig. 5. In the presence of the imperfection parameter $\epsilon$, we observe that the magnetization continues to exhibit oscillatory behavior with a period of $2T$, which is twice the intrinsic period of the Hamiltonian. As $\epsilon$ increases, the amplitude of the magnetization oscillations decreases, as indicated by the declining peak value in the FFT of the magnetization $m^{z}$ (see inset of Fig. 5). Despite this reduction, the magnetization oscillations remain stable over time. However, further increases in $\epsilon$ lead to a continued decline in oscillation amplitude, ultimately destabilizing the TC phase. For a chain length of $L=12$ (refer to Fig. 5), the critical value for stable magnetization oscillations is approximately $\epsilon_{c}\approx 0.3$; beyond this point, the oscillations become unstable and their amplitude diminishes. Notably, as the chain length increases, the stability of the TC phase improves, suggesting that the critical value of $\epsilon$ may rise for larger systems. Establishing the exact value of critical $\epsilon$ will necessitate further computational simulations on larger chains, which falls outside the scope of this study.
@@
Figure 5: The time evolution of the magnetization along the $z$-direction for various values of $\epsilon$ in a chain of length $L=12$ with coupling constant $J=5.5$ and external field $h=0.3$. The TC remains stable against a finite degree of imperfect rotation. The inset shows the FFT of the magnetization. As imperfection increases, the peak at $\omega=\pi$ diminishes and ultimately vanishes for larger $\epsilon$, signaling the breakdown of the TC phase.
@@ IV Time Quasicrystal (TQC): general definition
Let us consider a time-dependent Hamiltonian $\hat{H}(t)$ as followsr16 :
@@
When the frequencies $\omega_{j}$ in the Hamiltonian $H(t)$ are rationally related, the system exhibits periodic behavior, meaning that the dynamics of the system will repeat after a certain period $T$, which can be expressed as the least common multiple of the periods associated with the individual frequencies. In contrast, when the frequencies $\omega_{j}$ are irrationally related, the Hamiltonian becomes quasiperiodic.
Quasiperiodicity implies that the system does not repeat in a simple manner; instead, its state evolves over time in a way that covers a multi-dimensional toroidal space without returning to a previous configuration.
In the context of periodically and quasiperiodically driven systems, the concepts of time translation symmetry breaking (TTSB) and quasiperiodic time translation symmetry breaking (QTTSB) provide a framework for analyzing the behavior of operators over time.
For a periodically driven system with period $T$, TTSB occurs when the ensemble average of an operator $\hat{O}$ satisfies the condition $\langle\hat{O}(t+mT)\rangle=\langle\hat{O}(t)\rangle$ for integers $m\geq 2$. This indicates that the system exhibits a form of symmetry breaking where the observable properties of the system change over time in a different repeating manner.
On the other hand, in quasiperiodic systems, there is no single characteristic period due to the nature of their driving, which is typically defined by incommensurate frequencies. However, QTTSB can still be characterized through the frequency analysis of the system’s dynamics. Specifically, a quasiperiodic system demonstrates TQC behavior if there exists an operator $\hat{O}(t)$ such that its ensemble average can be expressed asr16 :
$\langle\hat{O}(t)\rangle=\sum_{\vec{n}}\hat{O}_{\vec{n}}\exp(i\sum_{i}\frac{n_{j}\omega_{j}}{p_{j}}t),$ (6)
where $\vec{n}$ are vectors of non-negative integers, $\omega_{j}$ are the driving frequencies, and $p_{j}$ are integers with at least one $p_{i}$ exceeding 1. This implies that the system exhibits a quasicrystalline ordering in time, characterized by multiple frequencies that do not lead to a simple periodic repetition but rather a complex structure reminiscent of quasicrystals in spatial dimensions.
In summary, TTSB refers to a breaking of periodic time symmetry in periodically driven systems, while QTTSB captures a similar notion for quasiperiodic systems through a more complex interplay of frequencies, indicating a richer dynamical structureFTS .
@@
The time-dependent Hamiltonian in Eq. (7) consists of three main terms: (1) ITF Hamiltonian $\hat{H}_{ITF}$ which describes the interactions in the system, including nearest-neighbor couplings and a transverse field. (2) Driving Hamiltonian $\hat{H}_{drive}$ which introduces a time-dependent driving force that oscillates sinusoidally. The driving terms involve both the $x$ and $y$-components of the spin operators, which can induce interesting dynamics in the system. (3) Additional magnetic field term $\frac{\Omega}{2}\sigma_{i}^{z}$, representing a static magnetic field in the $z$-direction, which can have a significant impact on the dynamics of the spins.
The time-dependent coefficients $h_{1}(t)$ and $h_{2}(t)$ are activated and deactivated periodically allowing for a controlled temporal evolution of the system. The definitions of $h_{1}(t)$ and $h_{2}(t)$ are critical for controlling the dynamics of the system. The piece wise definition ensures that the effects of the ITF and the driving terms do not overlap in time, allowing for distinct and potentially non-trivial time evolution dynamics. This can lead to a rich phase structure, especially when considering the interplay between disorder and driving.
We define $h_{1}(t)$ and $h_{2}(t)$ as follows:
$h_{1}(t)=\begin{cases}0,&\quad\frac{T}{4}>t\quad\text{and}\quad t<\frac{-T}{4}\\ 1,&\quad\frac{-T}{4}\leq t\leq\frac{T}{4}\end{cases}$ (10)
$h_{2}(t)=\begin{cases}\frac{1}{2},&\quad\frac{T}{4}>t\quad\text{and}\quad t<\frac{-T}{4}\\ 0,&\quad\frac{-T}{4}\leq t\leq\frac{T}{4}\end{cases}$
The schematic provided in FIG. 6 likely illustrates how the system transitions between the different Hamiltonian regimes over time. This visual representation can be very helpful for understanding the sequence of events that govern the system’s dynamics during the driving protocol.
Figure 6: An illustration of time intervals for the effects of various Hamiltonian terms.
The driving terms $h_{2}(t)\cos(\Omega t)$ and $h_{2}(t)\sin(\Omega t)$ in Eq. (9), are consistent with the definition of quasiperiodicity as specified in Eq. 5. This consistency can be confirmed through the Fourier expansion given by:
@@
where, $\omega=\frac{2\pi}{T}$ and $c_{n_{1}}$ are the coefficients of the expansion. The Fourier expansion of the driving terms indicates that they can exhibit quasiperiodic behavior, which is essential for the emergence of TQCs. The coefficients $c_{n_{1}}$ in the expansion will determine the specific nature of the quasiperiodicity, and the interplay between the various frequencies $n_{1}\omega+n_{2}\Omega$ can lead to complex dynamics.
@@ V TQC in the disordered ITF model
The phenomenon of QTTSB can be characterized through the temporal evolution of the magnetizations $m^{x,y,z}=\langle\hat{\sigma}^{x,y,z}\rangle$. While quasienergy level statistics are typically employed to investigate the thermalization phase transition in many-body systems, as noted in reference r8 , our quasiperiodically driven system presents challenges in defining the Floquet operator and quasienergies. Consequently, we focus on analyzing the time evolution of the magnetizations.
[…]
@@
The evolution of the magnetizations $m^{x,y,z}$ is initiated from the initial state $|\psi_{0}\rangle=\otimes_{i}^{L}[|\uparrow_{i}\rangle+|\downarrow_{i}\rangle]$. By averaging over 200 random realizations, we obtain the time dependence of the magnetizations $m^{x,y,z}$ over a time window of $100T$. The results, depicted in FIG. 7, reveal that when $\bar{J}\gg\bar{h}$, the subharmonic response is evident in the Fourier decomposition of the drive as expressed in Eq. (11). This response distinctly indicates QTTSB, characterized by sharp peaks at $k\frac{\omega}{2}\pm\Omega$ for odd integer $k$.
Figure 7: Top: Time evolution of the magnetizations $m^{x,y,z}$, for a chain of length $L=10$, with exchange coupling $J=5.5$ and transverse field $h=0.3$, revealing periodic oscillations in all three magnetization components. Bottom: FFT of the magnetization $m^{z}$ as a function of frequency $\omega$ showcasing distinct sharp peaks at $k\omega/2\pm\Omega$ (where $k$ is an odd integer), indicative of QTTSB and the formation of a QTC in the system.
@@
It is important to note that in certain systems, such as the Ising model under a longitudinal magnetic field, a transformation exists that can map a quasiperiodic Hamiltonian to a periodic one, enabling the transition from TQC to TC through a unitary transformation FTS . However, in the ITF model, such a transformation is not readily detectable. Specifically, the transformation must convert a quasiperiodic drive into a periodic one while leaving the time-independent Ising Hamiltonian unchanged. While this is achievable for the Ising model in a longitudinal field, identifying a similar transformation for the ITF chain poses significant challenges. As we will demonstrate in Section VI, although the ITF chain with couplings $J_{i}\in[-J/2,J/2]$ exhibits a TC phase under a periodic drive, it fails to exhibit a TQC phase under the quasiperiodic drive described in Eq. (9).

### v2

@@ III.1 Stability of the discrete TC phase under interactions
Figure 2: Time evolution of the magnetization along the $z$-direction for a chain of length $L=12$ subjected to an interaction $\lambda$ with varying strengths. The exchange couplings are randomly selected from the interval $[J/2,3J/2]$ with $J=5.5$, while the magnetic fields are drawn from the range [h,h][-h,h] with $h=0.3$. The envelope of the magnetization depends on the strength of the interaction $\lambda$. Increasing $\lambda$, leads to the instability of the oscillations with time period of $2T$, and consequently the destruction of the TC phase. FFT of the magnetization $m^{z}(t)$ exhibits a subharmonic peak at $\omega=\pi/2$. Increasing $\lambda$ results in a splitting of the Fourier peak, reflecting the beating of magnetization oscillations.
@@
In Fig. 2, we present the time evolution of the magnetization $m^{z}$ in the ferromagnetic many-body localized (FM-MBL) phase for varying interaction strengths. The observed time dependence of $m^{z}(t)$ reveals significant oscillatory behavior, characterized by a period of $2T$, superimposed on an oscillatory-decaying envelope that is dependent on the interaction strength $\lambda$. Specifically, for a finite-size system, the envelope exhibits a prolonged time period at low interaction strengths, and diminishes as $\lambda$ increases. Furthermore, we have computed the Fourier transform of $m^{z}(t)$ as a function of frequency for various $\lambda$ values. For small values of $\lambda$, within the observed time window $t/T\in[0,160]$, the magnetization oscillations remain smooth and exhibit no significant decay over time. The FFT analysis reveals a clear subharmonic peak at $\omega=\pi/2$. As $\lambda$ increases, this peak begins to split due to interference effects in the magnetization oscillations, eventually disappearing at a critical value $\lambda_{c}$, signaling the complete destruction of the TC phase. However, by extending the observation window, the decay of magnetization oscillations becomes more apparent over longer timescales, resulting in a double peak in the FFT even for small $\lambda$ values.
Figure 3: Time evolution of the magnetization along the $z$-direction for chains of varying lengths subjected to an interaction strength of $\lambda=1.5$. The exchange couplings are randomly selected from the interval $[J/2,3J/2]$ with $J=5.5$, while the magnetic fields are drawn from the range [h,h][-h,h] with $h=0.3$. As the system size increases, a significant enhancement in the stability of the magnetization oscillations is observed, characterized by a time period of $2T$. In the inset, the FFT of the magnetization $m^{z}(t)$ is presented. For smaller systems, such as with length $L=4$, the FFT, illustrated by the black symbols, reveals no sharp peak, indicating the absence of stable oscillations in $m^{z}(t)$. Conversely, as the system size increases, a pronounced peak emerges at the frequency $\omega=\pi/2$, signifying the establishment of stable oscillations in the magnetization.
@@
The observed oscillatory-decaying envelope in the magnetization can be attributed to finite-size effects. To substantiate this assertion, we present in Fig. 3 the time evolution of the magnetization $m^{z}(t)$ at a coupling strength $\lambda<J$ for chains of varying lengths. Notably, as the chain length increases, the period of the envelope oscillations also increases, suggesting that in the limit of infinite chain length ($L\to\infty$), the envelope disappears, leaving only stable oscillations with a period of $2T$. Furthermore, in the inset of Fig. 3, we present the FFT of $m^{z}(t)$ for different system lengths. For smaller systems, the FFT reveals a double peak structure, reflecting the beating of magnetization oscillations. As the system size increases, these double peaks converge into a single peak at frequency $\omega=\pi/2$. Continued increases in system size result in a sharper peak that becomes increasingly concentrated around $\omega=\pi/2$.
Figure 4: (Top: The absolute value of the magnetization $m^{z}(t)$ for a chain of length $L=12$ for different values of $\lambda$, with a coupling constant $J=5.5$ and an external field $h=0.3$. The TC remains stable for interaction strengths weaker than a critical value of approximately $0.4<\lambda_{c}<0.5$. Bottom: Finite-size scaling of the critical interaction strength $\lambda_{c}$ as a function of $1/L$, where $L$ denotes the system size. The data points represent numerical values, and the solid line corresponds to a linear fit. This extrapolation suggests that $\lambda_{c}$ approaches a finite, nonzero value as $L\to\infty$. Specifically, the thermodynamic limit value of the critical interaction strength is estimated to be $\lambda_{c}(\infty)\approx 0.67\pm 0.03$, indicating that the TC phase remains stable for interaction strengths below this threshold.
@@
We can evaluate the stability of the discrete TC phase against interactions by examining how the absolute value of the magnetization, $|m^{z}(t)|$, varies over time. As shown in Fig. 4-top, $|m^{z}(t)|$ remains constant over time for interaction strengths below a critical value $\lambda_{c}$, indicating that the TC phase is stable in this regime. The value of $\lambda_{c}$ depends on the system size; for a chain of length $L=12$, it lies approximately between 0.4 and 0.5.
To further understand the thermodynamic behavior, we performed a finite-size scaling analysis by plotting $\lambda_{c}$ as a function of $1/L$. As illustrated in Fig. 4-bottom, $\lambda_{c}$ increases with system size, and a linear fit suggests convergence to a finite, nonzero value in the thermodynamic limit. This strongly indicates that the TC phase persists over a finite range of interaction strengths and is not merely a finite-size artifact, confirming its robustness in the thermodynamic limit.
Refining the precise value of $\lambda_{c}$ necessitates analysis at larger system sizes. However, simulating chains of larger sizes presents substantial computational challenges, especially when a driving term complicates the dynamics. In this context, tensor network techniques, such as the time-evolving block decimation (TEBD) algorithm, offer a powerful solution. TEBD enables efficient representation and manipulation of quantum states, making it particularly well-suited for investigating quantum many-body systems. This approach is especially valuable when exact diagonalization becomes infeasible due to the exponential growth of the Hilbert space.
III.2 Stability of the discrete TC under imperfect rotations
@@
We now examine the stability of the TC observed in the ITF model under conditions of imperfect rotations. Instead of executing a perfect spin-flip rotation (i.e., a rotation around the $x$-axis by an angle of $g=\frac{\pi}{2}$), we implement an imperfect rotation characterized by an angle of $g=\frac{\pi}{2}-\epsilon$, where $\epsilon$ is a small positive value ($\epsilon<1$). Consequently, the time evolution operator is expressed as in Eq. (2) with the time parameters set to $t_{1}=t_{2}=1$.
@@
The time-dependent behavior of the magnetization along the $z$-direction is depicted in Fig. 5. In the presence of the imperfection parameter $\epsilon$, we observe that the magnetization continues to exhibit oscillatory behavior with a period of $2T$, which is twice the intrinsic period of the Hamiltonian. As $\epsilon$ increases, the amplitude of the magnetization oscillations decreases, as indicated by the declining peak value in the FFT of the magnetization $m^{z}(t)$ (see inset of Fig. 5). Despite this reduction, the magnetization oscillations remain stable over time. However, further increases in $\epsilon$ lead to a continued decline in oscillation amplitude, ultimately destabilizing the discrete TC phase. For a chain length of $L=12$ (refer to Fig. 5), the critical value for stable magnetization oscillations is approximately $\epsilon_{c}\approx 0.3$; beyond this point, the oscillations become unstable and their amplitude diminishes. Notably, as the chain length increases, the stability of the discrete TC phase improves. Establishing the exact value of critical $\epsilon$ will necessitate further computational simulations on larger chains, which falls outside the scope of this study.
@@
Figure 5: Time evolution of the magnetization along the $z$-direction for various values of $\epsilon$ in a chain of length $L=12$ with coupling constant $J=5.5$ and external field $h=0.3$. The discrete TC remains stable against a finite degree of imperfect rotation. The inset shows the FFT of the magnetization. As imperfection increases, the peak at $\omega=\pi/2$ diminishes and ultimately vanishes for larger $\epsilon$, signaling the disappearance of the discrete TC phase.
@@ IV Time Quasicrystal (TQC): general definition
Let us consider a time-dependent Hamiltonian $\hat{H}(t)$ as followsverdeny2016 :
@@
When the frequencies $\omega_{j}$ in the Hamiltonian $H(t)$ are rationally related, the system exhibits periodic behavior, meaning that the dynamics of the system will repeat after a certain period, which can be expressed as the least common multiple of the periods associated with the individual frequencies. In contrast, when the frequencies $\omega_{j}$ are irrationally related, the Hamiltonian becomes quasiperiodic. Quasiperiodicity implies that the system does not repeat in a simple manner; instead, its state evolves over time in a way that covers a multi-dimensional toroidal space without returning to a previous configuration.
In the context of periodically and quasiperiodically driven systems, the concepts of time translation symmetry breaking (TTSB) and quasiperiodic time translation symmetry breaking (QTTSB) provide a framework for analyzing the behavior of operators over time. For a periodically driven system with period $T$, TTSB occurs when the ensemble average of an operator $\hat{O}$ satisfies the condition $\langle\hat{O}(t+mT)\rangle=\langle\hat{O}(t)\rangle$ for integers $m\geq 2$. This indicates that the system exhibits a form of symmetry breaking where the observable properties of the system change over time in a different repeating manner.
On the other hand, in quasiperiodic systems, there is no single characteristic period due to the nature of their driving, which is typically defined by incommensurate frequencies. However, QTTSB can still be characterized through the frequency analysis of the system’s dynamics. Specifically, a quasiperiodic system demonstrates TQC behavior if there exists an operator $\hat{O}$ whose ensemble average satisfies the following Fourier transformation property verdeny2016 :
$\langle\hat{O}_{t}\rangle=\sum_{\vec{n}}\langle\hat{O}_{\vec{n}}\rangle\exp(i\vec{n}\cdot\vec{\tilde{\omega}}t),$ (7)
where $\vec{n}$ is a $d$-dimensional vector of integers, $\vec{\tilde{\omega}}$ is a $d$-dimensional vector with components $\tilde{\omega}_{1},\dots,\tilde{\omega}_{d}$, where $\tilde{\omega}_{j}=\omega_{j}/p_{j}$. Here, $\omega_{j}$ are the driving frequencies, and $p_{j}$ are non-negative integers with at least one $p_{i}$ exceeding 1. This implies that the system exhibits a quasicrystalline ordering in time, characterized by multiple frequencies that do not lead to a simple periodic repetition but rather a complex structure reminiscent of quasicrystals in spatial dimensionsFTS .
@@
The time-dependent Hamiltonian in Eq. (8) consists of three main terms:
(1) ITF Hamiltonian $\hat{H}_{ITF}$ which describes the interactions in the system, including nearest-neighbor couplings and a transverse field.
(2) Driving Hamiltonian $\hat{H}_{drive}$ which introduces a time-dependent driving force that oscillates sinusoidally. The driving terms involve both the $x$ and $y$-components of the spin operators, which can induce interesting dynamics in the system.
(3) Additional magnetic field term $\frac{\Omega}{2}\sigma_{i}^{z}$, representing a static magnetic field in the $z$-direction.
The time-dependent coefficients $h_{1}(t)$ and $h_{2}(t)$ are switched on and off periodically, enabling a controlled temporal modulation of the system’s evolution. Precisely defining $h_{1}(t)$ and $h_{2}(t)$ is essential for managing the system’s dynamics, as their piecewise structure prevents overlap between the effects of the ITF and the driving terms. This separation allows for a clear, stepwise evolution, which can give rise to non-trivial dynamical behavior.
We specify $h_{1}(t)$ and $h_{2}(t)$ as follows:
[…]
This construction ensures that, during the driving protocol, the system dynamically transitions through distinct Hamiltonian regimes in a well-defined sequence, thereby shaping its overall evolution.
According to Eq. (11), the driving terms $h_{2}(t)\cos(\Omega t)$ and $h_{2}(t)\sin(\Omega t)$ in Eq. (10), are consistent with the definition of quasiperiodicity as specified in Eq. 6. This consistency can be verified through their Fourier expansion, which is given by:
@@
where, $\omega=\frac{2\pi}{T}$ and $c_{n_{1}}$ are the coefficients of the expansion. The coefficients $c_{n_{1}}$ in the expansion will determine the specific nature of the quasiperiodicity, and the interplay between the various frequencies $n_{1}\omega+n_{2}\Omega$ can lead to complex dynamicsFTS .
@@ V TQC in the disordered ITF model
The phenomenon of QTTSB can be characterized through the temporal evolution of the magnetizations $m^{x,y,z}=\langle\hat{\sigma}^{x,y,z}\rangle$. While quasienergy level statistics are typically employed to investigate the thermalization phase transition in periodic systems, as noted in r8 , our quasiperiodic system presents challenges in defining the Floquet operator and quasienergies. In quasiperiodic systems, the absence of discrete TTS hinder the application of stroboscopic simplifications. Consequently, it is necessary to analyze the system’s dynamics through the time evolution operator $U(t)=\mathcal{T}\exp(-i\int_{0}^{t}H(t^{\prime})dt^{\prime})$, where $\mathcal{T}$ denotes time-ordering, and $H(t^{\prime})$ is the quasiperiodic Hamiltonian, given by Eq. (8).
For a chain of length $L=10$ with periodic boundary conditions, we employ exact diagonalization to compute the magnetization dynamics governed by the Hamiltonian presented in Eq. (8). The frequencies are fixed at $\omega=1$ ($T=2\pi$) and $\Omega=\sqrt{2}/4$, while the exchange couplings $J_{i}$ are randomly selected form the interval $[J/2,3J/2]$, and the transverse fields $h_{i}$ are drawn from the interval $[0,h]$. The magnetization data are sampled at a time step of $\frac{T}{200}$.
@@
The evolution of the magnetizations are initiated from the initial state $|\psi_{0}\rangle=\otimes_{i}^{L}[|\uparrow_{i}\rangle+|\downarrow_{i}\rangle]$. By averaging over 200 random realizations, we obtain the time dependence of the magnetizations over a time window of $100T$. The results, depicted in FIG. 6, reveal that when $J\gg h$ the subharmonic response is evident in the Fourier decomposition of the drive as expressed in Eq. (12). This response distinctly indicates QTTSB, characterized by sharp peaks at $k\frac{\omega}{2}\pm\Omega$ for odd integer $k$.
Figure 6: Top: Time evolution of the magnetization $m^{x}(t)$ for a chain of length $L=10$, with exchange coupling $J=5.5$ and transverse field $h=0.3$. The plot reveals periodic oscillations in $m^{x}(t)$, capturing the essential features of the quasiperiodic dynamics. The magnetization $m^{x}(t)$ exhibits an initial decay followed by the first plateau extending from $\tau_{\mathrm{pre}}$ to $\tau^{*}$. The persistence of coherent oscillations within this time interval highlights the stability of the prethermal regime, with its duration $t_{\mathrm{th}}=\tau^{*}-\tau_{\mathrm{pre}}$ defining the prethermal lifetime. Bottom: The FFT of $m^{x}(t)$ as a function of frequency $\omega$ exhibits sharp peaks at $k\omega/2\pm\Omega$, where $k$ is an odd integer. These peaks are indicative of QTTSB and signal the emergence of a TQC in the system.
The magnetization amplitude exhibits a smooth decline punctuated by temporal plateaus (see the envelope in Fig. 6). This behavior, which arises from the transverse field in the Ising chain, contrasts with that of a chain subjected to a longitudinal field, where the magnetization remains relatively stable over comparable timescales FTS . The observed decay of the envelope indicates that the stability of the TQC is limited; over sufficiently long times, the system heats up and approaches an infinite-temperature state dumitrescu2018 ; pre13 ; zhao2021 . This thermalization results from the quasiperiodic driving and occurs after a characteristic prethermal time, $t_{\mathrm{th}}$, which depends on the driving parameters dumitrescu2018 .
To quantify $t_{\mathrm{th}}$, we analyze the envelope of the magnetization $m^{x}(t)$. We identify two characteristic times from this envelope. The first, denoted $\tau_{\mathrm{pre}}$, marks the onset of the prethermal regime. It is defined as the time when the initial transient decay concludes and the magnetization enters its first nearly flat region (solid red line in Fig. 6). The second time, $\tau^{*}$, corresponds to the end of this first plateau, where the envelope begins its decay due to heating. We define the prethermal lifetime as $t_{\mathrm{th}}=\tau^{*}-\tau_{\mathrm{pre}}$, following the convention established in Ref. Kemp2017 ; prethermal2021 . This definition facilitates systematic comparisons across different system sizes and driving conditions, aligning our analysis with the broader literature on prethermal phases.
The first plateau is identified as the period during which the normalized magnetization $m^{x}(t)$ maintains an amplitude within the interval $0.9<m^{x}<1$.
@@
It is important to note that, in the Ising model subjected to a longitudinal magnetic field, a unitary transformation exists that maps the quasiperiodic Hamiltonian onto a periodic one. This transformation facilitates the transition from a TQC to a TC FTS . However, such a transformation does not exist for the ITF model. Specifically, the transformation would need to convert the quasiperiodic drive into a periodic one while leaving the static Ising Hamiltonian unchanged. While this is possible for the Ising model in a longitudinal field, no similar transformation is available for the ITF chain because the transverse term in the Hamiltonian does not commute with the interaction term.